# Parrondo’s paradox in susceptible-infectious-susceptible dynamics over periodic temporal networks

### v1

@@ Abstract
Many social and biological networks periodically change over time with daily, weekly, and other cycles. Thus motivated, we formulate and analyze susceptible-infectious-susceptible (SIS) epidemic models over temporal networks with periodic schedules. More specifically, we assume that the temporal network consists of a cycle of alternately used static networks, each with a given duration. We show Parrondo’s paradox behavior in this model, with which the periodic alternation of two static networks that are supra-threshold (i.e., above the epidemic threshold of the SIS model) yields a sub-threshold dynamics (i.e., with the number of infectious nodes exponentially decaying over time) in many cases. We find community structure to play an important role in shaping this phenomenon, and we study its dependence on the connectivity between and number of communities. Our findings associate such paradoxical behavior with anti-phase oscillatory behavior of the number of infectious nodes in different communities of the network.
@@ 1 Introduction
The connectivity of social contact networks is a main determinant for how contagions, i.e., spreading of communicative diseases, opinions, rumor, ideas, and so forth, occur in a population. Epidemic modeling on networks has been a useful tool that can help us understand, predict, and intervene into contagious processes occurring in the physical world and online [1, 2, 3, 4, 5, 6]. In fact, the structure of social and transportation networks can also dynamically change on the timescale of disease spreading. For example, human and animal mobility is a main obvious reason why contact networks vary over time. Airport transportation networks, which consist of airports as nodes and direct commercial flights between two airports as edges, also alter over days, weeks, and longer time [7, 8]. Time-varying networks are collectively called temporal networks, and epidemic processes on temporal networks have been extensively studied [9, 3, 10, 11, 12, 13, 14, 15, 16, 17]. Ignoring the time change of a network may lead to wrong or inaccurate conclusions about dynamics on networks. For example, the number of eventually infected nodes tends to be systematically different between simulations on a given temporal network and those on a counterpart static network that ignores temporal nature of the original network, even if the infection and recovery rates are assumed to be the same between the two sets of simulations (e.g., [12, 17]). Therefore, simulations on the static network would over-or under-estimate epidemic spread occurring on the given temporal network.
@@
One strategy for mathematically and computationally modeling temporal networks and dynamical processes on them is to use periodic temporal networks. Periodic temporal networks are useful because they allow us to derive and analyze a mapping of the given system’s state from one time point to one period after, reminiscent of the Poincaré map. A useful subclass of periodic temporal networks is periodic switching temporal networks, which are defined as time-varying networks in which one static network is used for a certain duration and then the network switches to another network, which is used for a certain duration, and so forth, and we come back to the first static network after the last one to complete one period, and then repeat the same cycle. It is common to use each static network for the same duration for simplicity. Periodic temporal networks in discrete time are by definition periodic switching temporal networks because each one static network is present in each time step, which generally changes over discrete times. Popularity of using periodic switching temporal networks is presumably due to their analytical tractability. Periodically switching networks have been applied to modeling, e.g., contagion processes [18, 19, 20, 21, 22, 23, 24] evolutionary dynamics [25, 26, 27, 28, 29, 30, 31, 32, 33], gene regulation [34, 35], network control [36, 37, 38, 39], and random walks [40, 41, 42].
@@
A phenomenon that can occur in periodically switching dynamical systems or stochastic processes is the Parrondo’s paradox. Originally, the Parrondo’s paradox refers to the situation in which a combination of two losing strategies generates a winning strategy [43, 44, 45]. When a person alternately and repetitively plays two different losing gambling games (i.e., losing regardless of the player’s strategy), the player may still win the game, even on average. The Parrondo’s paradox has also been found in theoretically motivated dynamical systems [46, 47, 48, 49, 50, 51, 52], social dynamics models [53, 54, 55, 56], evolutionary dynamics on networks [57], financial investment models [58, 59], chemical engineering [60], molecular transport [61], dynamics of allele frequency [62, 63, 64], cancer biology [65], cellular biology [66], and so on. Models of epidemic dynamics are no exception [67, 68, 69, 70, 71]. For example, Cheong et al. considered a model of COVID-19 outbreak [69]. In their model, when the community is not under lockdown, the infection rate is higher and the hospital cost is increased for individual persons. In contrast, when a lockdown is implemented, the total economic costs are increased despite the rate of epidemic spreading being reduced. They argue that implementing either one of these two strategies is a losing strategy. However, switching between the two strategies can reduce the overall cost per day; Cheong et al.  interpreted this to be a winning strategy, thereby representing a Parrondo’s paradox [69]. As another example, in [71], the authors implemented the Parrondo’s paradox idea on a periodic parameter switching algorithm with the aim of controlling the COVID-19 outbreak in their model.
@@
In the present article, we study the Parrondo’s paradox occurring on the susceptible-infectious-susceptible (SIS) dynamics over periodically switching temporal networks. We first find that the Parrondo’x paradox often occurs for small networks (i.e., those with a few nodes) modeling interaction between a small number of subpopulations of individuals. We corroborate our results with analysis of larger networks having the same structure of interaction between subpopulations. We characterize the observed paradoxical dynamics by quantifying anti-phase oscillation between the time courses of the fraction of infectious individuals in different subpopulations. We finally develop a perturbation theory, which also assists in explaining the reason for the paradoxical dynamics. Code used in this paper is available on GitHub [72].
@@ 2.1 SIS model over static networks
In this section, we present the SIS model over static networks and its epidemic threshold. We consider a static weighted network on $N$ nodes, $G(\mathcal{V},\mathcal{E})$, that may have self-loops; $\mathcal{V}=\{v_{1},\ldots,v_{N}\}$ is the set of nodes, and $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ is the set of edges. We denote the weighted adjacency matrix by $\textbf{A}=(a_{ij})$. By assumption, each node takes one of the two states, i.e., susceptible (denoted by S) or infectious (denoted by I), at any given time $t\in\mathbb{R}$. If node $v_{i}$ is susceptible and its neighbor $v_{j}$ is infectious, then $v_{j}$ infects $v_{i}$ at a constant infection rate, denoted by $\beta$ ($>0$). If $v_{i}$ is adjacent to multiple infectious nodes, each infectious neighbor infects $v_{i}$ independently of each other. In addition, any infectious node recovers to the susceptible state at a constant recovery rate, denoted by $\mu$ ($>0$). Therefore, the SIS model is a Markov process in continuous time with $2^{N}$ states.
The binary random variable, $X_{i}(t)$, encodes the state of $v_{i}$, i.e., $X_{i}(t)=0$ and $X_{i}(t)=1$ if $v_{i}$ is susceptible or infectious, respectively. Let $x_{i}(t)$ be the probability that the node $v_{i}$ is infectious at time $t$. Because there is no closed system of ODEs that describes the evolution of $x_{i}(t)$ (with $i\in\{1,\ldots,N\}$), a system of ODEs approximating the evolution of $x_{i}(t)$, called the individual-based-approximation (IBA), has been studied [5]. The IBA assumes independence of $\{x_{1}(t),\ldots,x_{N}(t)\}$. Under the IBA, the SIS dynamics is given by [73, 74, 75]
@@
where $X(t)=(X_{1}(t),\ldots,X_{N}(t))$, $\left\|X(0)\right\|_{1}=\sum_{i=1}^{N}X_{i}(0)$ represents the number of initially infectious nodes, and $\lambda_{\max}$ represents the largest eigenvalue of a matrix in modulus. Note that the Perron-Frobenius theorem guarantees that $\lambda_{\max}(\textbf{A})$ is real and positive, since the entries in A are nonnegative and at least one entry is positive.
@@ 2.3 SIS model over periodic temporal networks
Around the disease-free equilibrium, the linearized SIS dynamics (which assumes small $x_{i}(t)$ values) is given by
@@
To prove that $\lambda_{\max}(\mathcal{T})$ is real and positive, we show that $\mathcal{T}$ is a positive matrix. To this end, we start with
@@
where $\textbf{E}(T)=\int_{0}^{T}\textbf{A}(t)\text{d}t$. Because the identity matrix commutes with any matrix, $\textbf{E}(T)$ commutes with the identity matrix. Therefore, we obtain [79]
@@
Because $\beta>0$ and all the entries in $\textbf{A}(t)$ are nonnegative, all the entries of $\textbf{E}(T)$ are nonnegative. Let us assume that $\textbf{E}(T)$ is an irreducible matrix (i.e., $\left[e_{ij}(T)\right]^{k}>0$ $\forall i,j$ for some integer $k$). This condition is equivalent to assuming that the time-averaged network is strongly connected. The irreducibility of $\textbf{E}(T)$ implies that $e^{\beta\textbf{E}(T)}$ is a positive matrix (i.e., $e^{\beta e_{ij}(T)}>0$ $\forall i,j$) [79, 80]. This observation combined with Eq. (12) guarantees that $\mathcal{T}$ is a positive matrix. Therefore, $\lambda_{\max}(\mathcal{T})$ is real and positive owing to the Perron-Frobenius theorem.
The epidemic threshold is given by the value of $\beta$ at which $\lambda_{\max}(\mathcal{T})=1$. Eigenvalue $\lambda_{\max}(\mathcal{T})$ is equivalent to the largest Floquet multiplier in Floquet theory (see Appendix A for a brief review of a Floquet theory for periodic linear dynamical systems). We also define $\lambda_{\text{F}}\equiv T^{-1}\ln\lambda_{\max}(\mathcal{T})$, which gives the average growth rate of the infection. Under the IBA, the epidemic grows on a periodic temporal network if and only if
@@
thereby generalizing Eq. (6) for the temporal network case. Note that $\lambda_{\text{F}}>0$ is equivalent to $\lambda_{\max}(\mathcal{T})>1$ and vice versa. Because $\lambda_{\text{F}}$ is the largest of the Floquet exponents in Floquet theory (see Appendix A), we refer to $\lambda_{\text{F}}$ as the largest Floquet exponent in the following text.
@@
Note that $\mathcal{T}=e^{\beta\textbf{A}^{(\ell)}\tau_{\ell}}\cdots e^{\beta\textbf{A}^{(1)}\tau_{1}}e^{-\mu T}$ because any $e^{-\mu I\tau_{\ell^{\prime}}}$ $(=e^{-\mu\tau_{\ell^{\prime}}}I)$ commutes with any matrix and that each $e^{\beta\textbf{A}^{(\ell^{\prime})}\tau_{\ell^{\prime}}}$, $\ell^{\prime}\in\{1,\ldots,\ell\}$ is a positive matrix. Therefore, the Perron-Frobenius theorem guarantees that $\lambda_{\max}(\mathcal{T})$ is real and positive.
@@ 3 Results
In this section, we investigate the Parrondo’s paradox for the SIS dynamics over periodic switching networks. Specifically, epidemics on a temporal network can be sub-critical (i.e., the number of infectious nodes exponentially decays over time) over periods even if the SIS dynamics at any point of time is super-critical (i.e., the number of infectious nodes exponentially grows over time if the momentary static network is used forever). In section 3.1, we showcase the Parrondo’s paradox on a two-node periodic switching network. Section 3.2 shows the relationship between the amount of interaction and epidemic spreading. Section 3.3 illustrates the generality of the Parrondo’s paradox. We show relationships between the Parrondo’s paradox and anti-phase oscillation in section 3.4. Section 3.5 presents the perturbation theory for the largest Floquet exponent.
@@ 3.1 Parrondo’s paradox for a periodic switching network with two nodes
We start by studying the SIS dynamics on a minimal periodic switching network with two nodes and $\ell=2$ static networks to be switched between. We say that the Parrondo’s paradox occurs if the largest eigenvalues of ${\textbf{M}}^{(1)}\equiv\beta{\textbf{A}}^{(1)}-\mu{\textbf{I}}$ and that of ${\textbf{M}}^{(2)}\equiv\beta{\textbf{A}}^{(2)}-\mu{\textbf{I}}$, i.e., $\lambda_{\max}(\textbf{M}^{(1)})$ and $\lambda_{\max}(\textbf{M}^{(2)})$, are both positive and the largest Floquet exponent, $\lambda_{\text{F}}$, is negative. In this case, the SIS dynamics at any point of time is above the epidemic threshold such that the fraction of infectious nodes exponentially grows over time if either ${\textbf{A}}^{(1)}$ or ${\textbf{A}}^{(2)}$ is permanently used, whereas the epidemic exponentially decays over time on the periodic switching network. Generalizing this definition to the case of $\ell>2$ is straightforward. We only study the case of $\ell=2$ in this article.
Figure 2: Parrondo’s paradox for the periodic switching network with two nodes shown in Fig. 1. We set $\mu=0.5$ and vary $\beta$. (a) Largest Floquet exponent, $\lambda_{\text{F}}$ (black line), largest eigenvalue of $\textbf{M}^{(1)}$ (blue line), and that of $\textbf{M}^{(2)}$ (magenta line) as a function of $\beta$. The shaded region shows the range of $\beta$ in which the Parrondo’s paradox occurs. (b)–(e) Time courses of $x_{1}(t)$ and $x_{2}(t)$, which represent the infection probabilities in two nodes, for the four values of $\beta$ indicated in (a): (b) $\beta=0.01$, (c) $\beta=0.028$, (d) $\beta=0.032$, and (e) $\beta=0.04$.
@@
The shaded region in Fig. 2(a) represents the range of $\beta$ for which we have super-critical (i.e., above the epidemic threshold) dynamics on the static networks (i.e., $\lambda_{\max}(\textbf{M}^{(1)}),\lambda_{\max}(\textbf{M}^{(2)})>0$) and sub-critical dynamics on the periodic switching network (i.e., $\lambda_{\text{F}}<0$). In this situation, the Parrondo’s paradox is occurring such that the disease-free equilibrium is the unique stable equilibrium for Eq. (15), whereas, if the network were static then the disease-free equilibrium would be unstable and the number of infectious nodes would exponentially grow over time. Now we further explore the SIS dynamics at the $\beta$ values indicated by the circles in Fig. 2(a). In Fig. 2(b), (c), (d), and (e), we show $x_{1}(t)$ and $x_{2}(t)$ for $\beta=0.01$, $\beta=0.028$, $\beta=0.032$, and $\beta=0.04$ respectively. We recall that $x_{1}(t)$ and $x_{2}(t)$ are the probability that the $i$th node is infectious. We selected these four values of $\beta$ because each of them represents a qualitatively different situation regarding whether $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, or $\lambda_{\max}(\textbf{M}^{(2)})$ is positive or negative. In Fig. 2(b), $x_{1}(t)$ and $x_{2}(t)$ exponentially decay at any instantaneous time and in a long term. This result is expected because the growth rates of infection, $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$, are all negative at this value of $\beta$. The time courses of $x_{1}(t)$ and $x_{2}(t)$ show anti-phase behavior when $\beta$ is poised near (see Fig. 2(c)) or within (see Fig. 2(d)) the region in which the Parrondo’s paradox is present. In Fig. 2(e), both $x_{1}(t)$ and $x_{2}(t)$ exponentially grow over time in the long term. This last result is expected because $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$ are all positive at this value of $\beta$.
@@
To show that the Parrondo’s paradox can also be present in agent-based SIS dynamics, we simulated the stochastic, agent-based SIS model by the direct method of the Gillespie algorithm (e.g., [82]). To this end, we generated a network with $N=2000$ nodes and two equally sized communities using a stochastic block model (SBM) [83, 84]. See Appendix B for the network generation. We used the entries of $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$ to inform the edge probabilities for the periodic switching SBM and ran SIS simulations for the four values of $\beta$ used in Fig. 2(b)–(e). We assumed that there are initially 100 infectious nodes out of the 1000 nodes, which we selected uniformly at random, in each block, or community. The results shown in Appendix B indicate the presence of the Parrondo’s paradox in these stochastic simulations although the paradox’s outcome is less eminent than in the case of ODE simulations presented in this section. This discrepancy is partly due to the nonlinearity of the SIS dynamics, which is also present in the deterministic SIS dynamics, i.e., the IBA before the linearization (see Eq. (14)).

### v2

# Parrondo paradox in susceptible-infectious-susceptible dynamics over periodic temporal networks

@@ Abstract
Many social and biological networks periodically change over time with daily, weekly, and other cycles. Thus motivated, we formulate and analyze susceptible-infectious-susceptible (SIS) epidemic models over temporal networks with periodic schedules. More specifically, we assume that the temporal network consists of a cycle of alternately used static networks, each with a given duration. We observe a phenomenon in which two static networks are individually above the epidemic threshold but the alternating network composed of them renders the dynamics below the epidemic threshold, which we refer to as a Parrondo paradox for epidemics. We find that network structure plays an important role in shaping this phenomenon, and we study its dependence on the connectivity between and number of subpopulations in the network. We associate such paradoxical behavior with anti-phase oscillatory dynamics of the number of infectious individuals in different subpopulations.
@@ 1 Introduction
The connectivity of social contact networks is a main determinant for how contagions, i.e., spreading of communicative diseases, opinions, rumor, ideas, and so forth, occur in a population. Epidemic modeling on networks has been a useful tool that can help us understand, predict, and mitigate contagious processes occurring in the physical world and online [1, 2, 3, 4, 5, 6]. In fact, the structure of social and transportation networks can also dynamically change on the timescale of disease spreading. For example, human and animal mobility is a main obvious reason why contact networks vary over time. Airport transportation networks, which consist of airports as nodes and direct commercial flights between two airports as edges, also alter over days, weeks, and longer time [7, 8]. Time-varying networks are collectively called temporal networks, and epidemic processes on temporal networks have been extensively studied [9, 3, 10, 11, 12, 13, 14, 15, 16, 17]. Ignoring variations of a network over time may lead to wrong or inaccurate conclusions about dynamics on networks. For example, in the SIR process, the number of eventually infected nodes, called the final size, tends to be systematically different between simulations on a given temporal network and those on a counterpart static network (i.e., the aggregated, that is, time-averaged static network corresponding to the temporal network) that ignores temporal nature of the original network, even if the infection and recovery rates are assumed to be the same between the two sets of simulations (e.g., [12, 17]). Therefore, simulations on the static network would over-or under-estimate epidemic spread occurring on the given temporal network.
@@
One strategy for mathematically and computationally modeling temporal networks and dynamical processes on them is to use periodic temporal networks. Periodic temporal networks are useful because they allow us to derive and analyze a mapping of the given system’s state from one time point to one period after, reminiscent of the Poincaré map. A useful subclass of periodic temporal networks is periodic switching temporal networks, which are defined as time-varying networks in which one static network is used for a certain duration and then the network switches to another network, which is used for a certain duration, and so forth, and we come back to the first static network after the last one to complete one period, and then repeat the same cycle. It is common to use each static network for the same duration for simplicity. Periodic temporal networks in discrete time are by definition periodic switching temporal networks because each one static network is present in each time step, which generally changes over discrete times. Popularity of using periodic switching temporal networks is presumably due to their analytical tractability. Periodically switching networks have been applied to modeling, e.g., contagion processes [18, 19, 20, 21, 22, 23, 24] evolutionary dynamics [25, 26, 27, 28, 29, 30, 31, 32, 33], gene regulation [34, 35], network control [36, 37, 38, 39], and random walks [40, 41, 42, 43].
@@
A phenomenon that can occur in periodically switching dynamical systems or stochastic processes is the Parrondo paradox. Originally, the Parrondo paradox refers to the situation in which a combination of two losing strategies generates a winning strategy [44, 45, 46]. When a person alternately and repetitively plays two different losing gambling games (i.e., losing regardless of the player’s strategy), the player may still win the game, even on average. The Parrondo paradox has also been found in theoretically motivated dynamical systems [47, 48, 49, 50, 51, 52, 53], social dynamics models [54, 55, 56, 57], evolutionary dynamics on networks [58], financial investment models [59, 60], chemical engineering [61], molecular transport [62], dynamics of allele frequency [63, 64, 65], cancer biology [66], cellular biology [67], and so on. This phenomenon is also known in control theory community; a switching dynamical system composed of alternation of unstable dynamical systems can become stable [68, 69, 70, 71, 72, 73, 74]. Models of epidemic dynamics are no exception [75, 76, 77, 78, 79]. For example, Cheong et al. considered a model of COVID-19 outbreak [77]. In their model, when the community is not under lockdown, the infection rate is higher and the hospital cost is increased for individual persons. In contrast, when a lockdown is implemented, the total economic costs are increased despite the rate of epidemic spreading being reduced. They argue that implementing either one of these two strategies is a losing strategy. However, switching between the two strategies can reduce the overall cost per day; Cheong et al.  interpreted this to be a winning strategy, thereby representing a Parrondo’s paradox [77]. As another example, in [79], the authors implemented the Parrondo paradox idea on a periodic parameter switching algorithm with the aim of controlling the COVID-19 outbreak in their model. Thus, the literature suggests that the Parrondo paradox is useful for suppressing epidemic spreading.
@@
In the present study, we exploit this idea for potentially suppressing epidemic spreading over periodically switching networks with explicit network structure to investigate the effects of such structure. Specifically, we study the Parrondo paradox occurring on the susceptible-infectious-susceptible (SIS) dynamics over periodically switching temporal networks. We first find that the Parrondo paradox often occurs for small networks (i.e., those with a few nodes) modeling interaction between a small number of subpopulations of individuals. We corroborate our results with analysis of larger networks having the same structure of interaction between subpopulations. We characterize the observed paradoxical dynamics by quantifying anti-phase oscillation between the time courses of the fraction of infectious individuals in different subpopulations. We finally develop a perturbation theory, which also assists in explaining the reason for the paradoxical dynamics. Code used in this paper is available on GitHub [80].
@@ 2.1 SIS model over static networks
In this section, we present the SIS model over static networks and its epidemic threshold. We consider a static weighted network on $N$ nodes, $G(\mathcal{V},\mathcal{E})$, that may have self-loops; $\mathcal{V}=\{v_{1},\ldots,v_{N}\}$ is the set of nodes, and $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ is the set of edges. We denote the weighted adjacency matrix by $\textbf{A}=(a_{ij})$. By assumption, each node takes one of the two states, i.e., susceptible (denoted by S) or infectious (denoted by I), at any given time $t\in\mathbb{R}$. If node $v_{i}$ is susceptible and its neighbor $v_{j}$ is infectious, then $v_{j}$ infects $v_{i}$ at a constant rate $\beta a_{ij}$, where $\beta$ ($>0$) is the infection rate. If $v_{i}$ is adjacent to multiple infectious nodes, each infectious neighbor infects $v_{i}$ independently of each other. In addition, any infectious node recovers to the susceptible state at a constant recovery rate, denoted by $\mu$ ($>0$). Therefore, the SIS model is a Markov process in continuous time with $2^{N}$ states.
The binary random variable, $X_{i}(t)$, encodes the state of $v_{i}$, i.e., $X_{i}(t)=0$ and $X_{i}(t)=1$ if $v_{i}$ is susceptible or infectious, respectively. Let $x_{i}(t)$ be the probability that the node $v_{i}$ is infectious at time $t$. Because there is no closed system of ODEs that exactly describes the evolution of $x_{i}(t)$ (with $i\in\{1,\ldots,N\}$), a system of ODEs approximating the evolution of $x_{i}(t)$, called the individual-based-approximation (IBA), has been studied [5]. The IBA assumes independence of $\{x_{1}(t),\ldots,x_{N}(t)\}$. Under the IBA, the SIS dynamics is given by [81, 82, 83]
@@
where $X(t)=(X_{1}(t),\ldots,X_{N}(t))$, $\left\|X(0)\right\|_{1}=\sum_{i=1}^{N}X_{i}(0)$ represents the number of initially infectious nodes, and $\lambda_{\max}$ represents the largest eigenvalue of a matrix in modulus. Note that the Perron-Frobenius theorem guarantees that $\lambda_{\max}(\textbf{A})$ is real and positive; the theorem holds true because the entries in A are nonnegative and A is irreducible.
@@
Another interpretation of Eqs. (1), (2), (3), (4), and (6) is that a node represents a subpopulation of individuals [87, 88, 89]. In this case, the network is that of subpopulations, and $x_{i}(t)$ represents the fraction of infectious individuals in the $i$th subpopulation at time $t$. Figure 1 illustrates a network of individuals that are divided into two equal-sized subpopulations (with $10^{3}$ individuals each). The natural coarse graining of this network yields a $2\times 2$ matrix A, which encodes connectivity between the two subpopulations and within each subpopulation. To allow this interpretation, we allow the network to have self-loops (i.e., positive diagonal entries of A).
Figure 1: A network with $N=2000$ individuals organized in two subpopulations. (a) Adjacency matrix. Black dots indicate edges in the network, which are encoded by nonzero $A_{ij}$ entries. By grouping the first $10^{3}$ nodes into a subpopulation and the remaining $10^{3}$ nodes into another, we obtain a network composed of two subpopulations with self-loops. The subpopulation structure can be summarized by considering the probabilities of connection within and between subpopulations (which can be stored in a $2\times 2$ matrix). (b) Network visualization. Nodes (e.g., people) in the first and second subpopulation are colored blue and magenta, respectively.
@@ 2.3 SIS model over periodic temporal networks
Around the disease-free equilibrium, i.e., ${\bf x}(t)={\bf 0}$, the linearized SIS dynamics (which assumes small $x_{i}(t)$ values) is given by
@@
To prove that $\lambda_{\max}(\mathcal{T})$ is real and positive, we show that $\mathcal{T}$ is a matrix of which all the entries are positive under the assumption that the time-averaged network is strongly connected so that its adjacency matrix is nonnegative and irreducible. To this end, we start with
@@
where $\textbf{E}(T)=(e_{ij}(T))=\int_{0}^{T}\textbf{A}(t)\text{d}t$. (Note that one has ${\bf E}(T)=\sum_{\ell^{\prime}=1}^{\ell}\tau_{\ell^{\prime}}{\bf A}^{(\ell^{\prime})}$ for a periodic switching network.) Because the identity matrix commutes with any matrix, $\textbf{E}(T)$ commutes with the identity matrix. Therefore, we obtain [90]
@@
Because $\beta>0$ and all the entries in $\textbf{A}(t)$ are nonnegative, all the entries of $\textbf{E}(T)$ are nonnegative. Let us assume that $\textbf{E}(T)$ is an irreducible matrix (i.e., $[\textbf{E}(T)^{k}]_{ij}>0$ $\forall i,j$ for some integer $k$). This condition is equivalent to assuming that the time-averaged network is strongly connected. The irreducibility and nonnegativity of $\textbf{E}(T)$ implies that $e^{\beta\textbf{E}(T)}$ is a strictly positive matrix (i.e., $[e^{\beta\textbf{E}(T)}]_{ij}>0$ $\forall i,j$) [90, 91]. We show the proof in Appendix A. This observation, combined with Eq. (12) and $e^{-\mu T}>0$, guarantees that $\mathcal{T}$ a strictly positive matrix. Therefore, $\lambda_{\max}(\mathcal{T})$ is real and positive owing to the Perron-Frobenius theorem.
The epidemic threshold is given by the value of $\beta$ at which $\lambda_{\max}(\mathcal{T})=1$. Eigenvalue $\lambda_{\max}(\mathcal{T})$ is equivalent to the largest Floquet multiplier in Floquet theory (see Appendix B for a brief review of a Floquet theory for periodic linear dynamical systems). We also define $\lambda_{\text{F}}\equiv T^{-1}\ln\lambda_{\max}(\mathcal{T})$, which gives the average growth rate of the infection. Under the IBA, the epidemic grows on a periodic temporal network if and only if
@@
thereby generalizing Eq. (6) for the temporal network case. Note that $\lambda_{\text{F}}>0$ is equivalent to $\lambda_{\max}(\mathcal{T})>1$ and vice versa. Because $\lambda_{\text{F}}$ is the largest of the Floquet exponents in Floquet theory (see Appendix B), we refer to $\lambda_{\text{F}}$ as the largest Floquet exponent in the following text.
@@
Note that $\mathcal{T}=e^{\beta\textbf{A}^{(\ell)}\tau_{\ell}}\cdots e^{\beta\textbf{A}^{(1)}\tau_{1}}e^{-\mu T}$ because any $e^{-\mu I\tau_{\ell^{\prime}}}$ $(=e^{-\mu\tau_{\ell^{\prime}}}I)$ commutes with any matrix.
@@ 3 Results
In this section, we investigate the Parrondo paradox for the SIS dynamics over periodic switching networks. Specifically, epidemics on a temporal network can be sub-critical (i.e., the number of infectious nodes exponentially decays over time) over periods even if the SIS dynamics at any point of time is super-critical (i.e., the number of infectious nodes exponentially grows over time if the momentary static network is used forever). In section 3.1, we showcase the Parrondo paradox on a two-node periodic switching network. Section 3.2 shows the relationship between the amount of interaction and epidemic spreading. Section 3.3 illustrates the generality of the Parrondo paradox. We show relationships between the Parrondo paradox and anti-phase oscillation in section 3.4. Section 3.5 presents the perturbation theory for the largest Floquet exponent.
@@ 3.1 Parrondo’s paradox for a periodic switching network with two nodes
We start by studying the SIS dynamics on a minimal periodic switching network with two nodes and $\ell=2$ static networks to be switched between. We say that the Parrondo paradox occurs if the largest eigenvalues of ${\textbf{M}}^{(1)}\equiv\beta{\textbf{A}}^{(1)}-\mu{\textbf{I}}$ and that of ${\textbf{M}}^{(2)}\equiv\beta{\textbf{A}}^{(2)}-\mu{\textbf{I}}$, i.e., $\lambda_{\max}(\textbf{M}^{(1)})$ and $\lambda_{\max}(\textbf{M}^{(2)})$, are both positive and the largest Floquet exponent, $\lambda_{\text{F}}$, is negative. In this case, the SIS dynamics at any point of time is above the epidemic threshold such that the fraction of infectious nodes exponentially grows over time if either ${\textbf{A}}^{(1)}$ or ${\textbf{A}}^{(2)}$ is permanently used, whereas the epidemic exponentially decays over time on the periodic switching network. Generalizing this definition to the case of $\ell>2$ is straightforward. We only study the case of $\ell=2$ in this article.
Figure 3: Parrondo’s paradox for the periodic switching network with two nodes shown in Fig. 2. We set $\mu=0.5$ and vary $\beta$. (a) Largest Floquet exponent, $\lambda_{\text{F}}$ (black line), largest eigenvalue of $\textbf{M}^{(1)}$ (blue line), and that of $\textbf{M}^{(2)}$ (magenta line) as a function of $\beta$. The shaded region shows the range of $\beta$ in which the Parrondo paradox occurs. (b)–(e) Time courses of $x_{1}(t)$ and $x_{2}(t)$, which represent the infection probabilities in two nodes, for the four values of $\beta$ indicated in (a): (b) $\beta=0.01$, (c) $\beta=0.028$, (d) $\beta=0.032$, and (e) $\beta=0.04$.
@@
The shaded region in Fig. 3(a) represents the range of $\beta$ for which we have super-critical (i.e., above the epidemic threshold) dynamics on the static networks (i.e., $\lambda_{\max}(\textbf{M}^{(1)}),\lambda_{\max}(\textbf{M}^{(2)})>0$) and sub-critical dynamics on the periodic switching network (i.e., $\lambda_{\text{F}}<0$). In this situation, the Parrondo paradox is occurring such that the disease-free equilibrium is the unique stable equilibrium for Eq. (15), whereas, if the network were static then the disease-free equilibrium would be unstable and the number of infectious nodes would exponentially grow over time. Now we further explore the SIS dynamics at the $\beta$ values indicated by the circles in Fig. 3(a). In Fig. 3(b), (c), (d), and (e), we show $x_{1}(t)$ and $x_{2}(t)$ for $\beta=0.01$, $\beta=0.028$, $\beta=0.032$, and $\beta=0.04$ respectively. We recall that $x_{1}(t)$ and $x_{2}(t)$ are the probability that the $i$th node is infectious. We selected these four values of $\beta$ because each of them represents a qualitatively different situation regarding whether $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, or $\lambda_{\max}(\textbf{M}^{(2)})$ is positive or negative. In Fig. 3(b), $x_{1}(t)$ and $x_{2}(t)$ exponentially decay at any instantaneous time and in a long term. This result is expected because the growth rates of infection, $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$, are all negative at this value of $\beta$. The time courses of $x_{1}(t)$ and $x_{2}(t)$ show anti-phase behavior when $\beta$ is poised near (see Fig. 3(c)) or within (see Fig. 3(d)) the region in which the Parrondo paradox is present. In Fig. 3(e), both $x_{1}(t)$ and $x_{2}(t)$ exponentially grow over time in the long term. This last result is expected because $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$ are all positive at this value of $\beta$.
@@
To show that the Parrondo paradox can also be present in agent-based SIS dynamics, we simulated the stochastic, agent-based SIS model by the direct method of the Gillespie algorithm (e.g., [93]). To this end, we generated a network with $N=2000$ nodes and two equally sized communities using a stochastic block model (SBM) [94, 95]. See Appendix C for the network generation. We used the entries of $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$ to inform the edge probabilities for the periodic switching SBM and ran SIS simulations for the four values of $\beta$ used in Fig. 3(b)–(e). We assumed that there are initially 100 infectious nodes out of the 1000 nodes, which we selected uniformly at random, in each block, or community. The results shown in Appendix C indicate the presence of the Parrondo paradox in these stochastic simulations although the paradox’s outcome is less eminent than in the case of ODE simulations presented in this section. This discrepancy is partly due to the nonlinearity of the SIS dynamics, which is also present in the deterministic SIS dynamics, i.e., the IBA before the linearization (see Eq. (14)).